Solve the inequality:
|x+12/x-12|< or equal to 3
Please help

Answer by stanbon(75887)

Solve the inequality:
|x+12/x-12|<= 3
-------------
Looking you can see that x cannot be 12.
---
Solving the EQUALITY you gt:
(x+12)/(x-12) = 3 or (x+12)/(x-12) = -3
x+12 = 3x-26 or x+12 = -3x+36
2x = 38 or 4x = 24
x = 19 or x = 6
--------------------------------------
So, draw a number line and plot x = 6, x = 12, x = 19
--------------------------------------
Now find the solutions for the INEQUALITY by checking
a value in each of the four resulting intervals of
the number line.
--------------------------------------
|x+12/x-12| < 3
---
Let x = 0; you get |12/-12| < 3 ; true, so solutions in (-inf,6)
---
Let x = 10; you get |22/-2} < 3 ; false, so no solutions in (6,12)
---
Let x = 15; you get |27/3| < 3 ; false, so no solutions in (12,19)
---
Let x = 25 ; you get |37/13| < 3 ; true, so solutions in (19,+inf)
====================================================================
Combining the solutions for the EQUALITY and the INEQALITY you get:
(-inf,6] U [19,+inf)
-------------
Cheers,
Stan H.